OpenStudy (magbak):

Rewrite y = x2 + 14x + 29 in general form. (1 point) Rewrite y = 3x2 - 24x + 10 in general form. (1 point)

OpenStudy (anonymous):

y = x^2 + 14x + 29 is already in general form. In standard form, y = (x^2 + 14x + 49) - 20 = (x + 7)^2 - 20 y = 3x^2 - 24x + 20 = 3(x^2 - 8x + 16) - 38 = 3(x - 4)^2 - 38 (x - 9)^2 = 1 x - 9 = ± 1 x = 9 ± 1 x = 10 or x = 8

OpenStudy (magbak):

What is general form?